OpenStudy (anonymous):

what is the equation of the line parallel to the line y = 2x + 3 that passes through the point 1,4 in slope intercept form?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the slope of the given line

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

what is the equation of the given line

OpenStudy (anonymous):

y = 2x+3?

jimthompson5910 (jim_thompson5910):

what is the slope of that line

jimthompson5910 (jim_thompson5910):

hint: y = mx+b has a slope of m

OpenStudy (anonymous):

I have no idea.. do I input something?

jimthompson5910 (jim_thompson5910):

in y = 2x+3, what is the number in front of the x variable?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

so that is the slope

jimthompson5910 (jim_thompson5910):

any line parallel to y = 2x+3 will also have a slope of 2

jimthompson5910 (jim_thompson5910):

since parallel lines have equal slopes

OpenStudy (anonymous):

Right.. but how do I figure out the equation?

jimthompson5910 (jim_thompson5910):

how would you find the equation of the line if you knew the slope and a point that goes through it?

OpenStudy (anonymous):

I don't remember this at all

jimthompson5910 (jim_thompson5910):

we know that m = 2 we also know that x = 1 and y = 4 since we have the point (1,4) on this parallel line

jimthompson5910 (jim_thompson5910):

so... y = mx+b y = 2x + b ... plug in m = 2 4 = 2*1 + b ... plug in x = 1 and y = 4 solve for b

OpenStudy (anonymous):

Oh so you plug it in.. okay. 4 = b + 2

jimthompson5910 (jim_thompson5910):

which means b = ???

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

good, the equation of the parallel line is y = 2x + 2

OpenStudy (anonymous):

Can we do one other for practice?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

One sec.. so... What is the equation of the line perpendicular to the line y = 2x + 3 that passes through the point (1,4) in point-slope form? Would we use the same formula?

jimthompson5910 (jim_thompson5910):

it's a bit different with perpendicular lines

jimthompson5910 (jim_thompson5910):

the given slope is 2, which is really 2/1 flip the fraction: 2/1 -----> 1/2 flip the sign: 1/2 ----> -1/2

jimthompson5910 (jim_thompson5910):

so the perpendicular slope is -1/2

jimthompson5910 (jim_thompson5910):

so we know that m = -1/2 x = 1 y = 4 plug them all into y = mx + b then solve for b

OpenStudy (anonymous):

4 = -1/2(1) + b b = 9/2 would it be... y - 4 = -1/2(x-1)?

jimthompson5910 (jim_thompson5910):

you could use point slope form now get it into slope intercept form

jimthompson5910 (jim_thompson5910):

oh i see what you mean

jimthompson5910 (jim_thompson5910):

you got b = 9/2, so the equation in slope intercept form is \[\large y = -\frac{1}{2}x + \frac{9}{2}\]

jimthompson5910 (jim_thompson5910):

you could write it in point slope form to get y - 4 = -1/2(x-1)

OpenStudy (anonymous):

so I could use that solution?

jimthompson5910 (jim_thompson5910):

either one works, it will depend on what your book or teacher wants

OpenStudy (anonymous):

Yeah, I think the second one is what she wants. Thanks again. :)

jimthompson5910 (jim_thompson5910):

you're welcome